OpenStudy (anonymous):

simplify: 4t^2-1/ 15t+10 divided by 4t^2-4t+1/ 3t^2+5t+2 help please?

OpenStudy (turingtest):

the rule in general is\[\huge{\frac{a}{b}\over\frac{c}{d}}=\frac{a}{b}\cdot\frac{d}{c}\]

OpenStudy (turingtest):

so in your case we have\[\huge{\frac{4t^2-1}{15t+10}\over\frac{4t^2-4t+1}{3t^2+5t+2 }}=\frac{4t^2-1}{15t+10}\cdot\frac{3t^2+5t+2}{4t^2-4t+1}\]simplify completely and I bet some things will cancel

OpenStudy (anonymous):

so then what does that simplify to ?

OpenStudy (anonymous):

if these / means division all along then 4t^2 + t -1/15 + t 5/2t^2 +2

OpenStudy (anonymous):

\[\frac{4 t^2-1}{15 t+10}/\frac{4 t^2-4 t+1}{3 t^2+5 t+2}=\frac{2 t^2+3 t+1}{10 t-5}=\frac{t}{5}+\frac{2}{5}\text{ }\text{with } a \text{ remainder} \text{ of } 3 \]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You are welcome.